A particle is projected ...at that time its kinetic energy was KE1......and at the top point where vertical velocity is 0 its kinetic energy was KE2 .......


The ratio of KE2 : KE1 is 3:4 ........ find the difference in their velocities ......

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear ankit

let velocity at the begineeing ii u and angle of projection is θ.

KE1 =1/2 mu2

KE2 =1/2 m(ucosθ)2

given

KE2 : KE1 is 3:4

so cos2θ  =3/4

θ =30

difference in their velocity =u-ucosθ

                                     =u-u√3/2

                                     =u(1-√3/2)


    Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin